# thanks: Corresponding author

### v1

@@ Abstract
An accurate determination of singlet–triplet gaps in biradicals, including cyclobutadiene in the automerization barrier region where one has to balance the substantial nondynamical many-electron correlation effects characterizing the singlet ground state with the predominantly dynamical correlations of the lowest-energy triplet, remains a challenge for many quantum chemistry methods. High-level coupled-cluster (CC) approaches, such as the CC method with a full treatment of singly, doubly, and triply excited clusters (CCSDT), are often capable of providing reliable results, but the routine application of such methods is hindered by their high computational costs. We have recently proposed a practical alternative to converging the CCSDT energetics at small fractions of the computational effort, even when electron correlations become stronger and connected triply excited clusters are larger and nonperturbative, by merging the CC($P$;$Q$) moment expansions with the selected configuration interaction methodology abbreviated as CIPSI. We demonstrate that one can accurately approximate the highly accurate CCSDT potential surfaces characterizing the lowest singlet and triplet states of cyclobutadiene along the automerization coordinate and the gap between them using tiny fractions of triply excited cluster amplitudes identified with the help of relatively inexpensive CIPSI Hamiltonian diagonalizations.
@@ I Introduction
Biradicals play a key role in chemistry as reaction intermediates in thermal and photochemical pathwaysMyers et al. (1992); Pedersen et al. (1994); Berson (1978); Zgierski et al. (2005, 2005); Park et al. (2021); Abe et al. (2012) as well as functional materials used in molecular magnets,Dougherty (1991); Malrieu et al. (2014); Cho et al. (2016) battery electrodes,Sun et al. (2014) and organic photovoltaics.Minami and Nakano (2012); Hedley et al. (2017); Niklas and Poluektov (2017); Smith and Michl (2010) An important quantity characterizing the electronic structure of biradicals, especially in the context of designing molecules for magnetic, electrochemical, and photovoltaic applications, is the energy gap $\Delta E_{\text{S--T}}$ separating the lowest-lying singlet and triplet states (throughout this work, we define $\Delta E_{\text{S--T}}$ as $E_{\rm S}-E_{\rm T}$, where $E_{\rm S}$ and $E_{\rm T}$ are the electronic energies of the relevant singlet and triplet states, i.e., when the singlet is lower than the triplet, $\Delta E_{\text{S--T}}<0$). Accurate computational determination of the $\Delta E_{\text{S--T}}$ values in biradicals remains, however, a difficult task because it requires balancing the strong nondynamical many-electron correlation effects characterizing the low-spin singlet states with the predominantly dynamical correlations associated with their high-spin triplet counterparts. Slipchenko and Krylov (2002); Włoch et al. (2007); Li and Paldus (2008); Demel et al. (2008); Saito et al. (2011); Ess et al. (2011); Shen and Piecuch (2012); Abe (2013); Garza et al. (2014); Ibeji and Ghosh (2015); Ajala et al. (2017); Stoneburner et al. (2017); Zimmerman (2017); Shen and Piecuch (2021); Gulania et al. (2021); Boyn and Mazziotti (2021); Chakraborty et al. (2022)
@@
This challenge is exemplified by the cyclobutadiene molecule, which is the focus of the present study and which has fascinated experimental and theoretical chemists for decades with questions surrounding its low-lying electronic states, anti-aromaticity, and reactivity in cycloaddition and isomerization reactions. In the $D_{4\text{h}}$-symmetric square structure corresponding to the barrier along the automerization coordinate or the minimum on the lowest triplet potential, cyclobutadiene is a biradical with its four valence $\pi$ orbitals arranged in a network consisting of the nondegenerate $a_{2u}$ orbital, the doubly degenerate $e_{g}$ shell, and the nondegenerate $b_{1u}$ orbital. The distribution of two of the four valence electrons among the pair of degenerate frontier $e_{g}$ orbitals gives rise to three singlet states of the $B_{1g}(D_{4\text{h}})$, $A_{1g}(D_{4\text{h}})$, and $B_{2g}(D_{4\text{h}})$ symmetries and a $A_{2g}(D_{4\text{h}})$-symmetric triplet state, all involving the doubly occupied $a_{2u}$ and unoccupied $b_{1u}$ orbitals and the partially occupied $e_{g}$ shell in their zeroth-order description. Gulania et al. (2021); Craig (1950); Buenker and Peyerimhoff (1968); Nakamura et al. (1989); Balková and Bartlett (1994); Levchenko and Krylov (2004); Eckert-Maksić et al. (2006); Monino et al. (2022) As shown in the early ab initio calculations, Craig (1950); Buenker and Peyerimhoff (1968); Borden (1975); Kollmar and Staemmler (1977); Borden et al. (1978) and as confirmed in many subsequent theoretical studies, such as those reported in Refs. Gulania et al., 2021; Nakamura et al., 1989; Balková and Bartlett, 1994; Levchenko and Krylov, 2004; Eckert-Maksić et al., 2006; Saito et al., 2011; Ajala et al., 2017; Stoneburner et al., 2017; Zimmerman, 2017; Boyn and Mazziotti, 2021; Monino et al., 2022; Chakraborty et al., 2022, the lowest singlet of the $B_{1g}(D_{4\text{h}})$ symmetry, which has a substantial multiconfigurational character, is the ground state, whereas the predominantly single-reference $A_{2g}(D_{4\text{h}})$-symmetric triplet, in violation of Hund’s rule, is the first excited state. To be more specific, if one orients cyclobutadiene such that the two $C_{2}$ axes bisect the carbon–carbon bonds, which is a convention adopted in the present study, the ${}^{1}B_{1g}(D_{4\text{h}})$ ground state of the square structure is dominated by two closed-shell determinants in which one of the two degenerate $e_{g}$ orbitals is occupied by two electrons and the other one is empty (see, e.g., Refs. Borden et al., 1978; Nakamura et al., 1989; Levchenko and Krylov, 2004; Monino et al., 2022). This should be contrasted with the lowest-energy ${}^{3}A_{2g}(D_{4\text{h}})$ state, which is characterized by single occupancy of each of the $e_{g}$ orbitals.
@@
While the lowest ${}^{3}A_{2g}(D_{\text{4h}})$ state is stable in the square geometry, which is a minimum on the corresponding triplet surface, the ${}^{1}B_{1g}(D_{\text{4h}})$ ground state is unstable with respect to the rectangular distortion of the carbon–carbon bonds that lifts the degeneracy of the valence $e_{g}$ orbitals and lowers its total electronic energy due to the pseudo-Jahn–Teller effect. This results in the formation of the $D_{\text{2h}}$-symmetric rectangular species characterized by the closed-shell, predominantly single-determinantal, ${}^{1}A_{g}(D_{\text{2h}})$ ground state, which represents a minimum on the lowest-energy singlet potential. Gulania et al. (2021); Borden et al. (1978); Nakamura et al. (1989); Balková and Bartlett (1994); Levchenko and Krylov (2004); Boyn and Mazziotti (2021); Monino et al. (2022) The distortion of the multiconfigurational ${}^{1}B_{1g}(D_{4\text{h}})$ state into the ${}^{1}A_{g}(D_{2\text{h}})$ state coincides with the automerization coordinate in cyclobutadiene, which describes the conversion of the rectangular, $D_{2\text{h}}$-symmetric, closed-shell reactant (R) into the equivalent product conformer by passing through the square, $D_{4\text{h}}$-symmetric, biradical transition state (TS). Obtaining an accurate description of the potential energy curves (PECs) characterizing the lowest singlet [${}^{1}A_{g}(D_{2\text{h}})$] and triplet [${}^{3}B_{1g}(D_{2\text{h}})$] states of cyclobutadiene along its $D_{2\text{h}}$-symmetric automerization coordinate, and the gap between them, particularly in the neighborhood of the square biradical species, remains a significant challenge for modern ab initio techniques as it requires a high-level treatment of many-electron correlation effects in order to accurately capture and balance the strong nondynamical correlations associated with the multiconfigurational singlet state with the largely dynamical correlations dominating the triplet state.
@@
where $|\Phi\rangle$ is the $N$-electron reference determinant that serves as a Fermi vacuum and $T=\sum_{n=1}^{N}T_{n}$ is the cluster operator, with $T_{n}$ designating the $n$-body component of $T$ responsible for generating connected $n$-particle–$n$-hole ($n$p-$n$h) excitations out of $|\Phi\rangle$. It is well established that as long as the number of strongly correlated electrons is not too large, the standard hierarchy of CC approximations, including the CC method with singles and doubles (CCSD),Purvis and Bartlett (1982); Cullen and Zerner (1982); Scuseria et al. (1987); Piecuch and Paldus (1989) obtained by truncating $T$ at $T_{2}$, the CC approach with singles, doubles, and triples (CCSDT),Hoffmann and Schaefer (1986); Noga and Bartlett (1987); Scuseria and Schaefer (1988); Watts and Bartlett (1990) in which $T$ is truncated at $T_{3}$, the CC method with singles, doubles, triples, and quadruples (CCSDTQ),Oliphant and Adamowicz (1991); Kucharski and Bartlett (1991, 1992); Piecuch and Adamowicz (1994) where $T$ is truncated at $T_{4}$, and so on, rapidly converges to the exact, full configuration interaction (CI) limit. As a result, the single-reference CC approaches with a full treatment of higher-rank $T_{n}$ clusters with $n>2$, such as CCSDT or CCSDTQ, are often capable of accurately describing multireference situations, including substantial bond rearrangements in the course of chemical reactions and, what is especially important for this study, singlet–triplet gaps in biradical species, by capturing the relevant dynamical and nondynamical correlation effects via particle–hole excitations from a single determinant, without having to involve genuine multireference concepts.
@@
As shown in Figure 1, the lowest singlet and triplet PECs computed as functions of the $D_{\rm 2h}$-symmetric automerization coordinate $\lambda$ with full CCSDT are in very good agreement with their counterparts obtained with the CI method using perturbative selection made iteratively, Huron et al. (1973); Garniron et al. (2017, 2019) abbreviated as CIPSI, extrapolated to the full CI limit (see Section II for further details). The CCSDT energies are also very close to those determined using the double electron-attachment (DEA) equation-of-motion (EOM) CC methodologyMusiał et al. (2011, 2011); Shen and Piecuch (2013, 2014); Ajala et al. (2017); Gulania et al. (2021); Shen and Piecuch (2021) with a full treatment of 2p and 3p-1h and an active-space treatment of 4p-2h correlations on top of the CCSD description of the underlying closed-shell $({\rm C_{4}H_{4}})^{2+}$ core, Shen and Piecuch (2013, 2014); Ajala et al. (2017); Shen and Piecuch (2021) denoted as DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, where $N_{\text{u}}$ designates the number of active unoccupied orbitals of $({\rm C_{4}H_{4}})^{2+}$ included in the calculations to capture the leading 4p-2h effects in the target cyclobutadiene species (to accurately describe the 4p-2h effects associated with the valence orbitals of cyclobutadiene that correlate with the $e_{g}$ and $b_{1u}$ shells of the square TS structure, we set $N_{\text{u}}$ to 3). To illustrate the agreement between full CCSDT, perturbatively corrected and extrapolated CIPSI, and DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, which are three independent ab initio methodologies, we compare the vertical $\Delta E_{\text{S--T}}$ values at the R and TS geometries. When using the cc-pVDZDunning (1989) basis set, employed in the calculations reported in Figure 1 and the rest of this article, the CCSDT value of $\Delta E_{\text{S--T}}$ at the TS geometry is $-4.8$ kcal/mol. This is very close to $-5.2$ kcal/mol resulting from the state-of-the-art DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ calculations and $-5.9$ kcal/mol obtained with CIPSI. The CCSDT, DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, and perturbatively corrected and extrapolated CIPSI values of $\Delta E_{\text{S--T}}$ at the R geometry are $-30.6$, $-30.5$, and $-32.9$ kcal/mol, respectively, again in good agreement with one another.
@@
Given the high accuracy of the lowest singlet and triplet potential surfaces of cyclobutadiene and $\Delta E_{\text{S--T}}$ values along the $D_{\rm 2h}$-symmetric automerization pathway offered by full CCSDT, it may be tempting to turn to the approximate treatments of $T_{3}$ correlations that replace the expensive iterative $\mathscr{N}^{8}$ computational steps of CCSDT, where $\mathscr{N}$ is a measure of the system size, by the more practical $\mathscr{N}^{6}$ operations of CCSD combined with the noniterative $\mathscr{N}^{7}$ steps needed to correct the CCSD energetics for the leading $T_{3}$ correlations, as in the widely used CCSD(T) approach Raghavachari et al. (1989); Watts et al. (1993) or its more robust completely renormalized (CR) CC counterpart abbreviated as CR-CC(2,3).Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007) Unfortunately, neither CCSD(T) and CR-CC(2,3) nor any of the other noniterative triples corrections to CCSD, such as CCSD(T)Λ,Stanton (1997); Crawford and Stanton (1998); Kucharski and Bartlett (1998) CCSD(2)T,Hirata et al. (2001, 2004); Gwaltney and Head-Gordon (2000, 2001) CR-CCSD(T)Piecuch and Kowalski (2000); Kowalski and Piecuch (2000); Piecuch et al. (2002, 2004) and its locally renormalized extension,Kowalski and Piecuch (2005) $\Lambda$-CCSD(T),Taube and Bartlett (2008, 2008) and CCSD(${\rm T}\mbox{--}n$),Eriksen et al. (2014, 2014) are capable of providing accurate results when the coupling of the lower-order $T_{1}$ and $T_{2}$ components of the cluster operator with their higher-rank $T_{3}$ counterpart becomes large. For example, even the most robust triples correction to CCSD defining CR-CC(2,3), which improves CCSD(T) and other similar approaches in situations involving electronic quasi-degeneracies, such as those present in single bond breaking, Shen and Piecuch (2012); Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007); Piecuch et al. (2007); Ge et al. (2007, 2008) struggles in describing the PEC of the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the neighborhood of the automerization barrier region, where $T_{3}$ clusters become large, nonperturbative, and strongly coupled to $T_{1}$ and $T_{2}$.Shen and Piecuch (2012) As a result, as shown in Ref. Chakraborty et al., 2022 and this study, the CR-CC(2,3) $\Delta E_{\text{S--T}}$ value at the singlet TS structure, of 4.4 kcal/mol when the cc-pVDZ basis is used, is in large error (including incorrect sign) relative to its CCSDT $-4.8$ kcal/mol counterpart [CCSD(T) gives 3.9 kcal/mol, which is similarly inaccurate].

### v2

@@ Abstract
ABSTRACT: An accurate determination of singlet–triplet gaps in biradicals, including cyclobutadiene in the automerization barrier region where one has to balance the substantial nondynamical many-electron correlation effects characterizing the singlet ground state with the predominantly dynamical correlations of the lowest-energy triplet, remains a challenge for many quantum chemistry methods. High-level coupled-cluster (CC) approaches, such as the CC method with a full treatment of singly, doubly, and triply excited clusters (CCSDT), are often capable of providing reliable results, but the routine application of such methods is hindered by their high computational costs. We have recently proposed a practical alternative to converging the CCSDT energetics at small fractions of the computational effort, even when electron correlations become stronger and connected triply excited clusters are larger and nonperturbative, by merging the CC($P$;$Q$) moment expansions with the selected configuration interaction methodology abbreviated as CIPSI. We demonstrate that one can accurately approximate the highly accurate CCSDT potential surfaces characterizing the lowest singlet and triplet states of cyclobutadiene along the automerization coordinate and the gap between them using tiny fractions of triply excited cluster amplitudes identified with the help of relatively inexpensive CIPSI Hamiltonian diagonalizations.
@@ I Introduction
Biradicals play a key role in chemistry as reaction intermediates in thermal and photochemical pathwaysMyers et al. (1992); Pedersen et al. (1994); Berson (1978); Zgierski et al. (2005, 2005); Park et al. (2021); Abe et al. (2012) as well as functional materials used in molecular magnets,Dougherty (1991); Malrieu et al. (2014); Cho et al. (2016) battery electrodes,Sun et al. (2014) and organic photovoltaics.Minami and Nakano (2012); Hedley et al. (2017); Niklas and Poluektov (2017); Smith and Michl (2010) An important quantity characterizing the electronic structure of biradicals, especially in the context of designing molecules for magnetic, electrochemical, and photovoltaic applications, is the energy gap $\Delta E_{\text{S--T}}$ separating the lowest-lying singlet and triplet states (throughout this work, we define $\Delta E_{\text{S--T}}$ as $E_{\rm S}-E_{\rm T}$, where $E_{\rm S}$ and $E_{\rm T}$ are the electronic energies of the relevant singlet and triplet states, i.e., when the singlet is lower than the triplet, $\Delta E_{\text{S--T}}<0$). Accurate computational determination of the $\Delta E_{\text{S--T}}$ values in biradicals remains, however, a difficult task because it requires balancing the strong nondynamical many-electron correlation effects characterizing the low-spin singlet states with the predominantly dynamical correlations associated with their high-spin triplet counterparts. Slipchenko and Krylov (2002); Włoch et al. (2007); Li and Paldus (2008); Demel et al. (2008); Saito et al. (2011); Ess et al. (2011); Shen and Piecuch (2012); Abe (2013); Garza et al. (2014); Ibeji and Ghosh (2015); Ajala et al. (2017); Stoneburner et al. (2017); Zimmerman (2017); Shen and Piecuch (2021); Gulania et al. (2021); Boyn and Mazziotti (2021); Chakraborty et al. (2022); Demel et al. (2023); Szabados et al. (2025)
@@
This challenge is exemplified by the cyclobutadiene molecule, which is the focus of the present study and which has fascinated experimental and theoretical chemists for decades with questions surrounding its low-lying electronic states, anti-aromaticity, and reactivity in cycloaddition and isomerization reactions. In the $D_{4\text{h}}$-symmetric square structure corresponding to the barrier along the automerization coordinate or the minimum on the lowest triplet potential, cyclobutadiene is a biradical with its four valence $\pi$ orbitals arranged in a network consisting of the nondegenerate $a_{2u}$ orbital, the doubly degenerate $e_{g}$ shell, and the nondegenerate $b_{1u}$ orbital. The distribution of two of the four valence electrons among the pair of degenerate frontier $e_{g}$ orbitals gives rise to three singlet states of the $B_{1g}(D_{4\text{h}})$, $A_{1g}(D_{4\text{h}})$, and $B_{2g}(D_{4\text{h}})$ symmetries and a $A_{2g}(D_{4\text{h}})$-symmetric triplet state, all involving the doubly occupied $a_{2u}$ and unoccupied $b_{1u}$ orbitals and the partially occupied $e_{g}$ shell in their zeroth-order description. Gulania et al. (2021); Craig (1950); Buenker and Peyerimhoff (1968); Nakamura et al. (1989); Balková and Bartlett (1994); Levchenko and Krylov (2004); Eckert-Maksić et al. (2006); Monino et al. (2022) As shown in the early ab initio calculations, Craig (1950); Buenker and Peyerimhoff (1968); Borden (1975); Kollmar and Staemmler (1977); Borden et al. (1978) and as confirmed in many subsequent theoretical studies, such as those reported in Refs. Gulania et al., 2021; Nakamura et al., 1989; Balková and Bartlett, 1994; Levchenko and Krylov, 2004; Eckert-Maksić et al., 2006; Saito et al., 2011; Ajala et al., 2017; Stoneburner et al., 2017; Zimmerman, 2017; Boyn and Mazziotti, 2021; Monino et al., 2022; Chakraborty et al., 2022; Demel et al., 2023; Szabados et al., 2025, the lowest singlet of the $B_{1g}(D_{4\text{h}})$ symmetry, which has a substantial multiconfigurational character, is the ground state, whereas the predominantly single-reference $A_{2g}(D_{4\text{h}})$-symmetric triplet, in violation of Hund’s rule, is the first excited state. To be more specific, if one orients cyclobutadiene such that the two $C_{2}$ axes bisect the carbon–carbon bonds, which is a convention adopted in the present study, the $^{1}B_{1g}(D_{4\text{h}})$ ground state of the square structure is dominated by two closed-shell determinants in which one of the two degenerate $e_{g}$ orbitals is occupied by two electrons and the other one is empty (see, e.g., Refs. Borden et al., 1978; Nakamura et al., 1989; Levchenko and Krylov, 2004; Monino et al., 2022). This should be contrasted with the lowest-energy $^{3}A_{2g}(D_{4\text{h}})$ state, which is characterized by single occupancy of each of the $e_{g}$ orbitals.
@@
While the lowest $^{3}A_{2g}(D_{\text{4h}})$ state is stable in the square geometry, which is a minimum on the corresponding triplet surface, the $^{1}B_{1g}(D_{\text{4h}})$ ground state is unstable with respect to the rectangular distortion of the carbon–carbon bonds that lifts the degeneracy of the valence $e_{g}$ orbitals and lowers its total electronic energy due to the pseudo-Jahn–Teller effect. This results in the formation of the $D_{\text{2h}}$-symmetric rectangular species characterized by the closed-shell, predominantly single-determinantal, $^{1}A_{g}(D_{\text{2h}})$ ground state, which represents a minimum on the lowest-energy singlet potential. Gulania et al. (2021); Borden et al. (1978); Nakamura et al. (1989); Balková and Bartlett (1994); Levchenko and Krylov (2004); Boyn and Mazziotti (2021); Demel et al. (2023); Szabados et al. (2025); Monino et al. (2022) The distortion of the multiconfigurational $^{1}B_{1g}(D_{4\text{h}})$ state into the $^{1}A_{g}(D_{2\text{h}})$ state coincides with the automerization coordinate in cyclobutadiene, which describes the conversion of the rectangular, $D_{2\text{h}}$-symmetric, closed-shell reactant (R) into the equivalent product conformer by passing through the square, $D_{4\text{h}}$-symmetric, biradical transition state (TS). Obtaining an accurate description of the potential energy curves (PECs) characterizing the lowest singlet $^{1}A_{g}(D_{2\text{h}})$] and triplet $^{3}B_{1g}(D_{2\text{h}})$] states of cyclobutadiene along its $D_{2\text{h}}$-symmetric automerization coordinate, and the gap between them, particularly in the neighborhood of the square biradical species, remains a significant challenge for modern ab initio techniques as it requires a high-level treatment of many-electron correlation effects in order to accurately capture and balance the strong nondynamical correlations associated with the multiconfigurational singlet state with the largely dynamical correlations dominating the triplet state.
@@
where $|\Phi\rangle$ is the $N$-electron reference determinant that serves as a Fermi vacuum and $T=\sum_{n=1}^{N}T_{n}$ is the cluster operator, with $T_{n}$ designating the $n$-body component of $T$ responsible for generating connected $n$-particle–$n$-hole (nnp-nnh) excitations out of $|\Phi\rangle$. It is well established that as long as the number of strongly correlated electrons is not too large, the standard hierarchy of CC approximations, including the CC method with singles and doubles (CCSD),Purvis and Bartlett (1982); Cullen and Zerner (1982); Scuseria et al. (1987); Piecuch and Paldus (1989) obtained by truncating $T$ at $T_{2}$, the CC approach with singles, doubles, and triples (CCSDT),Hoffmann and Schaefer (1986); Noga and Bartlett (1987); Scuseria and Schaefer (1988); Watts and Bartlett (1990) in which $T$ is truncated at $T_{3}$, the CC method with singles, doubles, triples, and quadruples (CCSDTQ),Oliphant and Adamowicz (1991); Kucharski and Bartlett (1991, 1992); Piecuch and Adamowicz (1994) where $T$ is truncated at $T_{4}$, and so on, rapidly converges to the exact, full configuration interaction (CI) limit. As a result, the single-reference CC approaches with a full treatment of higher-rank $T_{n}$ clusters with $n>2$, such as CCSDT or CCSDTQ, are often capable of accurately describing multireference situations, including substantial bond rearrangements in the course of chemical reactions and, what is especially important for this study, singlet–triplet gaps in biradical species, by capturing the relevant dynamical and nondynamical correlation effects via particle–hole excitations from a single determinant, without having to involve genuine multireference concepts.
@@
[…]
@@
Given the high accuracy of the lowest singlet and triplet potential surfaces of cyclobutadiene and $\Delta E_{\text{S--T}}$ values along the $D_{\rm 2h}$-symmetric automerization pathway offered by full CCSDT, it may be tempting to turn to the approximate treatments of $T_{3}$ correlations that replace the expensive iterative $\mathscr{N}^{8}$ computational steps of CCSDT, where $\mathscr{N}$ is a measure of the system size, by the more practical $\mathscr{N}^{6}$ operations of CCSD combined with the noniterative $\mathscr{N}^{7}$ steps needed to correct the CCSD energetics for the leading $T_{3}$ correlations, as in the widely used CCSD(T) approach Raghavachari et al. (1989); Watts et al. (1993) or its more robust completely renormalized (CR) CC counterpart abbreviated as CR-CC(2,3).Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007) Unfortunately, neither CCSD(T) and CR-CC(2,3) nor any of the other noniterative triples corrections to CCSD, such as CCSD(T)Λ,Stanton (1997); Crawford and Stanton (1998); Kucharski and Bartlett (1998) CCSD(2)T,Hirata et al. (2001, 2004); Gwaltney and Head-Gordon (2000, 2001) CR-CCSD(T)Piecuch and Kowalski (2000); Kowalski and Piecuch (2000); Piecuch et al. (2002, 2004) and its locally renormalized extension,Kowalski and Piecuch (2005) $\Lambda$-CCSD(T),Taube and Bartlett (2008, 2008) and CCSD(${\rm T}\mbox{--}n$),Eriksen et al. (2014, 2014) are capable of providing accurate results when the coupling of the lower-order $T_{1}$ and $T_{2}$ components of the cluster operator with their higher-rank $T_{3}$ counterpart becomes large. For example, even the most robust triples correction to CCSD defining CR-CC(2,3), which improves CCSD(T) and other similar approaches in situations involving electronic quasi-degeneracies, such as those present in single bond breaking, Shen and Piecuch (2012); Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007); Piecuch et al. (2007); Ge et al. (2007, 2008) struggles in describing the PEC of the lowest $^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the neighborhood of the automerization barrier region, where $T_{3}$ clusters become large, nonperturbative, and strongly coupled to $T_{1}$ and $T_{2}$.Shen and Piecuch (2012) As a result, as shown in Ref. Chakraborty et al., 2022 and this study, the CR-CC(2,3) $\Delta E_{\text{S--T}}$ value at the singlet TS structure, of 4.4 kcal/mol when the cc-pVDZ basis is used, is in large error (including incorrect sign) relative to its CCSDT 4.8-4.8 kcal/mol counterpart. CCSD(T) gives 3.9 kcal/mol, which is similarly inaccurate. This is because neither CR-CC(2,3) nor CCSD(T), nor any other noniterative triples correction to CCSD, can correctly describe $T_{3}$ contributions to $\Delta E_{\text{S--T}}$ that at the square TS geometry of cyclobutadiene are a few times larger, in absolute value, than the CCSDT singlet–triplet gap itself. For example, the $T_{3}$ effects on $\Delta E_{\text{S--T}}$ at the TS structure, estimated by subtracting the singlet–triplet gap obtained in the CCSD/cc-pVDZ calculations from its CCSDT/cc-pVDZ counterpart, are 15.1-15.1 kcal/mol, i.e., they are more than three times larger than the value of $\Delta E_{\text{S--T}}$ obtained in the CCSDT/cc-pVDZ calculations. The fact that the $T_{3}$ contributions to the singlet–triplet gap of cyclobutadiene in the vicinity of the automerization barrier region become so massive and difficult to describe by the noniterative triples corrections to CCSD is closely related to the dramatic increase of $T_{1}$, $T_{2}$, and $T_{3}$ cluster amplitudes characterizing the lowest-energy $^{1}A_{g}(D_{2\text{h}})$ state, especially the amplitudes defining $T_{2}$ and $T_{3}$ operators that engage valence orbitals around the Fermi level, reflecting on the increasingly strongly correlated character of this state as one transitions from the R to TS structures. For example, the largest $T_{1}$, $T_{2}$, and $T_{3}$ amplitudes obtained in the CCSDT/cc-pVDZ calculations for the $A_{g}(D_{2\text{h}})$-symmetric singlet ground state of cyclobutadiene at its R geometry are 0.032682-0.032682, 0.206225-0.206225, and $0.003414$, respectively. At the TS geometry, they become 0.054928-0.054928, 0.895785-0.895785, and $0.014836$, respectively, i.e., we observe a four-fold increase in the largest $T_{2}$ and $T_{3}$ amplitudes compared to the R structure. This should be contrasted with the behavior of the lowest-energy triplet state, where the largest $T_{1}$, $T_{2}$, and $T_{3}$ cluster amplitudes resulting from the CCSDT/cc-pVDZ computations remain relatively small and barely change when the ${\rm R}\rightarrow{\rm TS}$ ($\lambda=0\rightarrow 1$) geometrical transformation is examined. They are $0.031220$ for $T_{1}$, 0.163848-0.163848 for $T_{2}$, and $0.003285$ for $T_{3}$ at $\lambda=0$ and 0.030476, 0.137065-0.137065, and $0.002555$, respectively, when $\lambda$ becomes 1. It is this drastically different behavior of the lowest $^{1}A_{g}(D_{2\text{h}})$ and $^{3}B_{1g}(D_{2\text{h}})$ states of cyclobutadiene and the rapid growth of the cluster amplitudes characterizing the former state that engage valence orbitals, especially those associated with $T_{2}$ and $T_{3}$, as the automerization barrier region is approached, which result in failures of methods such as CCSD(T) and CR-CC(2,3) in describing the corresponding singlet–triplet gap.